what mass of nitric oxide canbe made from the reaction of 8.00g NH3 with 17.0g O2
1
Expert's answer
2020-10-09T13:44:15-0400

STEP I: BALANCED EQUATION

4NH3(g) + 5O2(g) "\to" 4NO(g) + 6H2O(l)


STEP II: MOLES OF EACH REAGENT PRESENT

Moles of ammonia


"Moles =\dfrac{mass}{MM}"


Molar mass of NH3 is ( N=14, H=1)

MM = 14 +3 = 17g/mol



"Moles =\dfrac{8g}{17g} = 0.471 mol"


Moles of oxygen (MM of O2 = 2 x 16 = 32g/mol)



"Moles =\dfrac{17g}{32g/mol} = 0.25mol"


STEP III: DETERMINE THE LIMITING REAGENT

Moles of oxygen required to react with 0.472mol of ammonia


Mole ratios from the equation (NH3: O2) = 4:5


Thus moles of O2 = "\dfrac{5}{4}x0.471mol" = 0.589mol.

But there are only 0.25 mol.

Therefore oxygen is the limiting reagent

The amount of product depends on amount of oxygen present.


STEP IV: DETERMINE AMOUNT OF PRODUCT

Moles of NO produced


Mole ratios of O2:NO = 5:4


5 mole of O2 produces 4 moles of NO

0.25 moles of O2 would produce?


Moles of NO "=\dfrac{4}{5}x 0.25mol = 0.2 mol"


Mass of NO = Moles x MM (MM = 14 +16)


Mass of NO = 0.2 mol x 30g/mol


= 6g on NO